OpenStudy (anonymous):

Give an equation of the line through the point (6, 5) and parallel to the line 2x + 3y = 7.

OpenStudy (anonymous):

You should know one thing that Parallel Lines have same Slope..

OpenStudy (anonymous):

Here, the question says, that find the equation of a line (say l) which passes through a point (x,y) and parallel to another line (say m) with equation 2x + 3y = 7.. So, as Slopes of both the parallel lines will be same, so first step is to find the slope of the line m having equation : 2x + 3y = 7

OpenStudy (anonymous):

Can you find the slope of equation : 2x + 3y = 7 ??

OpenStudy (anonymous):

-2/3?

OpenStudy (anonymous):

Oh, you are real fast..

OpenStudy (anonymous):

This is slope of line m, so, line l should have same slope as both the lines are parallel..

OpenStudy (anonymous):

So, line l has slope of -2/3 and you know you are given that it is passing through a point.. So, use point slope form and find the equation now.. Can you do that??

OpenStudy (anonymous):

y-5=-2/3(x-6)

OpenStudy (anonymous):

Point Slope Form is given as: \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

Yeah just simplify it, so that it looks nice while seeing it.. :)

OpenStudy (anonymous):

Good, you are going real good.. :) Multiply by 3 and bring the x term on Left Hand Side...

OpenStudy (anonymous):

OK i came up with y=-2/3x-1 is that riht?

OpenStudy (anonymous):

*right

OpenStudy (anonymous):

Check it once again..

OpenStudy (anonymous):

y-5=-2/3x+4

OpenStudy (anonymous):

Better is to write it in the form in which the other line m is given to you: ie : \(a_x + by = c\)

OpenStudy (anonymous):

*\(ax\)

OpenStudy (anonymous):

Multiply by 3 both the sides first and tell me what did you get??

OpenStudy (anonymous):

ok so i came up with y=-2/3x+9 oh wait where did you get the 3?

OpenStudy (anonymous):

Look here: You got : \[y - 5 = -\frac{2}{3}(x -6)\] Here multiply by 3 both the sides and now tell me what have you got?

OpenStudy (anonymous):

that gives me y-15=-2x-12 ?

OpenStudy (anonymous):

-12 or +12??

OpenStudy (anonymous):

Be careful with signs..

OpenStudy (anonymous):

ohhh ummmm +12 sorry

OpenStudy (anonymous):

y-12=-2x+12

OpenStudy (anonymous):

which in turn gives me y=-2x+24?

OpenStudy (anonymous):

Are you alright?? Have you studied more than your capacity??

OpenStudy (anonymous):

Oh Yes I'm fine XD I'm sorry what did I do wrong?

OpenStudy (anonymous):

I have some trouble with these types of problems

OpenStudy (anonymous):

On LHS, how you got y - 12??

OpenStudy (anonymous):

You have applied method just superbly, just you are doing mistakes in calculation. Go slow and be careful..

OpenStudy (anonymous):

*headdesk* I have no idea how I got that... Let me redo that starting with y-5=-2/3(x-6)

OpenStudy (anonymous):

Yes, now multiply 3 both the sides first.. Just do what I am saying to you..

OpenStudy (anonymous):

Ok sorry i'm confused what do i multiply by 3?

OpenStudy (anonymous):

Wait, I should have explain it earlier to you.. Your line m is in which form??

OpenStudy (anonymous):

m=-2/3

OpenStudy (anonymous):

2x + 3y = 7, the form : \(ax + by = c\) this is called Standard Form..

OpenStudy (anonymous):

No doubt, you have got y-5 = -2/3(x-6) but it does not seem like ax + by = c to me.. Does it look like to you???

OpenStudy (anonymous):

oHHH OK so i need to switch y-5=-2/3(x-6) to standard form?

OpenStudy (anonymous):

Yes, this is what I was trying to say.. :)

OpenStudy (anonymous):

Oh Ok sorry! that gives me 2x+3y=0

OpenStudy (anonymous):

Do whatever way you know, I want to bring just x and y on LHS and constants on RHS.. Now do in your way..

OpenStudy (anonymous):

NO..

OpenStudy (anonymous):

Oh sorry... Ok you said times it by 3 what do i times by 3?

OpenStudy (anonymous):

There is a reason behind it..

OpenStudy (anonymous):

If you don't want to do it, then do like this: \[y - 5 = -\frac{2}{3}(x - 6)\] Multiply -2/3 inside brackets: \[y - 5 = -\frac{2x}{3} + 4\] Now bring x on LHS, so add 2x/3 \[y - 5 + \frac{2x}{3} = 4 \implies \frac{2x}{3} + y = 9\] Got it till here?

OpenStudy (anonymous):

OK so 2/3x+y=9 now what?

OpenStudy (anonymous):

You will take LCM here or what...

OpenStudy (anonymous):

Or here, also time it with 3, so you will easily remove 3 from denominator.. My main purpose is to remove denominator terms, denominator of x and y terms should be 1..

OpenStudy (anonymous):

OHHH I think I got it now! Thanks!

OpenStudy (anonymous):

You sure??

OpenStudy (anonymous):

Finally you get : \(2x + 3y = 27\)

OpenStudy (anonymous):

Thanks!